OpenStudy (anonymous):

idodine molecules dissociates into atoms after absorbing light of 4500 ANGSTROM .if one quantum of radiation is absorbed by each molecules ,calculate the kinetic energy of iodine atoms .

OpenStudy (unklerhaukus):

Energy is conserved Energy of the photon is \[E_\gamma=h\nu=\frac{hc}{\lambda}\] Combined Kinetic Energy of Both iodine atoms \[KE_{I_2}=\frac{mv^2}{2}\] Kinetic energy of each atom\[KE_I=\frac{mv^2}{4}\]